# Bi-Lipschitz embedding metric triangles in the plane

### v1

@@ Abstract
footnotetext: 2020 Mathematics Subject Classification. 51F30, 30L05. M. Romney was partially supported by the National Science Foundation under grant DMS-2413156. The main results of this paper form a part of the Senior Research Project of X. Luo and A. L. Tao.
1 Introduction
The bi-Lipschitz embedding problem asks one to given an intrinsic characterization of those metric spaces that can be embedded in some Euclidean space under a bi-Lipschitz map. Moreover, if a metric space does embed in some Euclidean space, one may then attempt to find or estimate the optimal target dimension and bi-Lipschitz distortion. The bi-Lipschitz embedding problem is wide open in full generality and has produced a rich body of research from a variety of mathematical points of view. We refer the reader to [5], [6], [7, Chapter 15], [8], [10] and [11] for general background.
In this article, we study a specific class of metric spaces, namely, metric polygons, and their embeddings in Euclidean spaces, with a focus on metric triangles. We say that a metric polygon is a metric space comprised of finitely many closed intervals (i.e., metric spaces each isometric to $[0,a]$ for some $a>0$), called edges, joined in cyclic fashion. For two points lying on different edges, there is no constraint on the distance between them beyond what follows from the triangle inequality. A metric polygon with three edges is called a metric triangle, with four edges is called a metric quadrilateral, and so forth. See 2.4 below for a formal definition.
To our knowledge, metric polygons first appear explicitly in the recent paper [9], in which it is shown that any metric triangle embeds in $\mathbb{R}^{2}$ (equipped with the Euclidean metric) under a bi-Lipschitz map with distortion at most $12$. The fact is then applied to the problem of finding good polyhedral approximations of metric surfaces under minimal geometric assumptions. We refer to the embedding found in [9] as the tripodal embedding; it will be reviewed in Section 3 below. Our goal is to explore the topic of embedding metric polygons more systematically and extend the embedding result from [9] in several directions.
@@ 1 Introduction
In Section 4.4, we use an example that we call the twisted heart to show that the constant $4\sqrt{7/3}$ cannot be improved. This leaves open the question of the best possible value for the distortion $D$ with an arbitrary embedding. Let $\operatorname{lip}(\triangle)$ denote the minimum value for which any metric triangle embeds in $\mathbb{R}^{2}$ with distortion at most $\operatorname{lip}(\triangle)$. A simple example, called the three-petal rose, shows that $\operatorname{lip}(\triangle)\geq 2$; see Section 4.2. The example of the twisted heart suggests an even larger value for $\operatorname{lip}(\triangle)$, perhaps roughly the value $4$. However, we have not been able to find any general embedding method that improves upon the tripodal embedding.
@@
In Section 4, we study the previously mentioned examples, along with two other representative examples: tripods and the intrinsic circle. Specifically, we consider the problem of finding distortion-minimizing bi-Lipschitz embeddings for these triangles. We can think of such a distortion-minimizing embedding as an optimal “Euclidean realization” of the original triangle. Note, however, that such an embedding is typically not unique; since the distortion of an embedding depends only on the worst pair of points in the space, one can typically perturb a given embedding elsewhere without affecting the distortion.
In the final main section this paper, Section 5, we consider the situation for arbitrary metric polygons. Note that it follows immediately from general properties of bi-Lipschitz embeddings (see Theorem 3.2 in [6]) that every metric $n$-gon bi-Lipschitz embeds in $\mathbb{R}^{2n+1}$ with uniformly bounded distortion depending only on $n$. Thus our main question is, for a given $n$, to determine the smallest dimension for which one obtains an embedding of every $n$-gon with uniformly bounded distortion. We show that the analogue of 1.1 is false for metric quadrilaterals.
@@
There is a sequence of metric quadrilaterals $X_{n}$ such that $\operatorname{lip}(X_{n},\mathbb{R}^{2})$ is unbounded as $n\to\infty$.
@@
We also expect that every metric quadrilateral is individually bi-Lipschitz embeddable in $\mathbb{R}^{2}$, i.e., without a uniform bound on the bi-Lipschitz constant.
@@
On the other hand, the two fundamental examples of non-planar graphs—$K_{5}$, the complete graph on five vertices, and $K_{3,3}$, the complete bipartite graph on two sets of three vertices—can each be realized as a metric pentagon. See 5.3 below.
In addition to our work, there have been other investigations into the metric geometry of classes of one-dimensional metric spaces. For example, bi-Lipschitz embeddings of so-called quasiconformal trees have also been studied in [3] and [4].
Acknowledgement
We thank Guy C. David for feedback on a draft of our paper.
2 Background
2.1 Metric spaces and bi-Lipschitz maps
@@
Finally, the concatenation of two curves $\Gamma_{1}\colon[a,b]\to X$, $\Gamma_{2}\colon[c,d]\to X$ satisfying $\Gamma_{1}(b)=\Gamma_{2}(c)$ is denoted by $\Gamma_{1}*\Gamma_{2}$. This is the curve from $[a,b+d-c]$ to $X$ defined by the formula $\Gamma_{1}(t)$ if $t\leq b$ and $\Gamma_{2}(t-b+c)$ if $t\geq b$.
2.2 Metric polygons
@@
Let $n\geq 2$. A metric $n$-gon is a metric space that is the union of $n$ subspaces $I_{1},\ldots,I_{n}$ that are each isometric to a closed interval $[0,\ell_{n}]$ and connected at the endpoints cyclically. More precisely, if we let $\varphi_{j}\colon[0,\ell_{j}]\to I_{j}$ denote the required isometry onto the interval $I_{j}$, then it holds that $\varphi_{j}(\ell_{j})=\varphi_{j+1}(0)$ for all $j\in\{1,\ldots,n\}$, with the identification $n+1=1$. A metric $3$-gon is called a metric triangle, and a metric $4$-gon is called a metric quadrilateral, and so forth. More generally, a metric polygon is a metric $n$-gon for some $n$.
@@
Note that this definition allows the sets $I_{j}$ to intersect at points other than the endpoints, although the nature of such additional intersection points is constrained by the triangle inequality. Also, note that the terms “edge” and “vertex” here are distinct from the topological meaning of these terms for a graph. For example, graphs (such as $K_{5}$ or $K_{3,3}$, as discussed in 5.3) can be represented as metric polygons. In these cases, the topological edges and vertices of the graph need not coincide with its edges and vertices as a metric polygon.
Next, we introduce some main examples of metric triangles that will be studied further in Section 4. These are illustrated in Figure 1. In each of these examples, the space is a length space or subset of a length space.
@@
    To view this space as a metric triangle, we pick any three points that do not all belong to the same semicircle to be vertices. For example, we can take $\theta_{1}=0$, $\theta_{2}=2\pi/3$ and $\theta_{3}=4\pi/3$.
@@
    The three-petal rose. Let $S_{1},S_{2},S_{3}$ be three copies of the intrinsic circle as just defined. Let $R$ be the wedge sum of $S_{1},S_{2},S_{3}$ with basepoint 00 in each $S_{i}$ (that is, the quotient space of the disjoint union of $S_{1},S_{2},S_{3}$ given by identifying the point 00 in each $S_{i}$). We define $d$ to be the length metric on $R$. More precisely, $d$ is defined as follows. If $\theta,\tau\in S_{i}$ for the same $i$, then $d(\theta,\tau)$ agrees with the metric in (a). If $\theta\in S_{i}$ and $\tau\in S_{j}$ for $i\neq j$, then $d(\theta,\tau)=d(\theta,0)+d(0,\tau)$.
    To view this space as a metric triangle, we take the antipodal point on each circle as vertices. That is, let $p=\pi\in S_{1}$, $q=\pi\in S_{2}$ and $r=\pi\in S_{3}$. The points 00 and $p$ divide $S_{1}$ into two semicircles $S_{1}^{1}$, $S_{1}^{2}$. Likewise, $S_{2}$ and $S_{3}$ are divided into semicircles $S_{2}^{1},S_{2}^{2}$ and $S_{3}^{1},S_{3}^{2}$, respectively. Then the edges of $R$ are $[pq]=S_{1}^{2}\cup S_{2}^{1}$, $[qr]=S_{2}^{2}\cup S_{3}^{1}$ and $[rp]=S_{3}^{2}\cup S_{1}^{1}$.
@@
    Tripods. Choose values $l_{i}\geq 0$, $i=1,2,3$, and consider the three intervals $T_{i}=[0,l_{i}]$, $i=1,2,3$. We define $T$ to be the wedge sum of $T_{1},T_{2},T_{3}$ with basepoint 00 in each $T_{i}$. We define $d$ to be the length metric on $T$. That is, if $x,y\in T_{i}$ for the same $i$, then $d(x,y)=|x-y|$. If $x\in T_{i}$ and $y\in T_{j}$ for $i\neq j$, then $d(x,y)=d(x,0)+d(0,y)=x+y$. Such a metric space is called a tripod, or a metric tripod for emphasis. We call the point 00 the center of the tripod.
    Let $p=l_{1}\in T_{1}$, $q=l_{2}\in T_{2}$, $r=l_{3}\in T_{3}$. We can view $T$ as a metric triangle by taking the points $p,q,r$ as vertices and $[pq]=T_{1}\cup T_{2}$, $[qr]=T_{2}\cup T_{3}$, $[rp]=T_{3}\cup T_{1}$ as edges.
@@
    The twisted heart. The example is less obvious but turns out to be the extremal example for the tripodal embedding. We start with the graph consisting of fourteen edges of unit length as shown in Figure 1(d). The distance between two points is the length of the shortest curve between them. We obtain a triangle by taking $p,q,r$ as vertices and removing the edges from $a$ to $c$ and from $b$ to $e$ (the red edges in Figure 1(d)) without changing the metric on the remaining points. Observe that each pair of vertices is at distance $4$ apart. Moreover, $d(a,b)=4$ while $d(a,c)=d(b,e)=1$. We denote this example by $H$.
$p$$q$$r$
$p$$q$$r$
$p$$q$$r$
$p$$q$$r$$a$$b$$e$$c$
Figure 1: Examples of metric polygons.
Our final definition, used in the proof of 1.1, is the Gromov product, which measures the deficit in the triangle inequality for a triple of points.
@@
Let $\triangle$ be a metric triangle with vertices $p,q,r$. Let $T$ be the tripod formed as in 2.5 by taking $l_{1}=(q\cdot r)_{p}$, $l_{2}=(p\cdot r)_{q}$ and $l_{3}=(p\cdot q)_{r}$. Define a map $P\colon\triangle\to T$ by setting $P(p)=l_{1}\in T_{1}$, $P(q)=l_{2}\in T_{2}$, $P(r)=l_{3}\in T_{3}$, and mapping each edge of $\triangle$ by arc length onto the corresponding edge of $T$. Then $P$ is $1$-Lipschitz.
@@
3 Proof of optimal distortion bound
In this section, we give the proof of our main result, 1.1, giving the optimal distortion bound for the tripodal embedding of an arbitrary metric triangle.
We start by reviewing the definition of the tripodal embedding as given in [9]. Let $\triangle$ be an arbitrary metric triangle with edges $I_{1},I_{2},I_{3}$. Denote the metric on $\triangle$ by $d$. Let $p$ be the common vertex of $I_{1}$ and $I_{2}$, $q$ be the common vertex of $I_{2}$ and $I_{3}$, and $r$ be the common vertex of $I_{3}$ and $I_{1}$. For each edge $I_{j}$, let $\widehat{I}_{j}$ denote the union of the other two edges of $\triangle$. We can associate to the triangle $\triangle$ a metric tripod $T$ with matching edge lengths and corresponding map $P\colon\triangle\to T$ as in 2.7. To this $T$ we associate an embedded tripod $Y$ and map $\varphi\colon T\to Y$ following the notation of the previous section. For conciseness, we write $\bar{x}$ in place of $(\varphi\circ P)(x)$. Let $v_{1}=(\frac{1}{2},\frac{\sqrt{3}}{2})$, $v_{2}=(-1,0)$ and $v_{3}=(\frac{1}{2},-\frac{\sqrt{3}}{2})$. Then $F\colon\triangle\to\mathbb{R}^{2}$ is defined by the formula
@@ 3 Proof of optimal distortion bound
$p$$q$$r$$o_{1}$$o_{2}$$o_{3}$$y$$x$
$p$$q$$r$$o_{1}$$o_{2}$$o_{3}$$x$$y$
$p$$q$$r$$o_{1}$$o_{2}$$o_{3}$$x$$y$
$p$$q$$r$$o_{1}$$o_{2}$$o_{3}$$x$$y$
$p$$q$$r$$o_{1}$$o_{2}$$o_{3}$$x$$y$
Figure 2: The five cases in the proof of 1.1.
@@
We fix some additional notation. Let $I_{1}^{p}$ denote the subarc of $I_{1}$ consisting of those points $x\in I_{1}$ for which $d(x,p)\leq(q\cdot r)_{p}$. Define $I_{1}^{q}$, $I_{2}^{q}$, $I_{2}^{r}$, $I_{3}^{r}$ and $I_{3}^{p}$ similarly. For each edge $I_{j}$, let $o_{j}$ denote the unique point such that $\bar{o}_{j}$ is the origin. Moreover, we let $z\in\triangle$ denote a point satisfying $d(x,z)=\operatorname{dist}(x,\widehat{I}_{1})$. This point may fail to be uniquely determined, so we fix a choice of $z$. Similarly, we let $w\in\triangle$ denote a point satisfying $d(y,w)=\operatorname{dist}(y,\widehat{I}_{j})$, where $j$ is such that $y\in I_{j}$.
In this proof, we consider two arbitrary distinct points $x,y\in\triangle$ and show that $\|F(x)-F(y)\|$ satisfies appropriate bounds. By symmetry, we may consider five individual cases in our proof of 1.1: (1) $x,y\in I_{1}^{p}$; (2) $x\in I_{1}^{p},y\in I_{1}^{q}$; (3) $x\in I_{1}^{p},y\in I_{3}^{p}$; (4) $x\in I_{1}^{p},y\in I_{3}^{r}$; (5) $x\in I_{1}^{q},y\in I_{3}^{r}$. See Figure 2. Each of these cases is represented by one of the lemmas in this section.
@@
Assume that $d(x,p)<d(y,p)$. We first verify the lower bound. Note that both $\bar{x}$ and $\bar{y}$ lie on the positive real axis, with $\|\bar{x}-\bar{y}\|=d(x,y)$. Moreover, $F(x)$ and $F(y)$ lie on the lines $\{\bar{x}+tv_{1}:t\in\mathbb{R}\}$ and $\{\bar{y}+tv_{1}:t\in\mathbb{R}\}$, respectively. The distance between these lines is $\frac{\sqrt{3}}{2}d(x,y)$. See Figure 3(a). This verifies the lower bound.
Next, we show the upper bound. Observe that $|d(x,\widehat{I}_{1})-d(y,\widehat{I}_{2})|\leq d(x,y)$. Write $F(y)$ as $F(y)=\bar{y}+t(y)v_{1}$ for some $t(y)\geq 0$. Then $F(x)$ belongs to the set $\bar{x}+tv_{j}$, where $t\in[t(y)-d(x,y),t(y)+d(x,y)]$. The point in this set furthest from $F(y)$ is $\bar{x}+(t(y)+d(x,y))v_{j}$, which is at distance $\sqrt{3}d(x,y)$ from $F(y)$. See Figure 3(b). This verifies the upper bound. ∎
@@
Assume without loss of generality that $d(o_{1},y)\leq d(o_{1},x)$. We first verify the lower bound. As in the previous case, observe that $F(x)$ and $F(y)$ lie on the lines $\{\bar{x}+tv_{1}:t\in\mathbb{R}\}$ and $\{\bar{y}+tv_{1}:t\in\mathbb{R}\}$, respectively. The distance between these lines is $\frac{\sqrt{3}}{2}d(o_{1},x)+\frac{\sqrt{3}}{2}d(o_{1},y)$. See Figure 4(a). Since $d(x,y)=d(x,o_{1})+d(o_{1},y)$, we obtain the lower bound.
We next verify the upper bound. Write $F(y)$ as $F(y)=\bar{y}+t(y)v_{1}$ for some $t(y)\geq 0$. As in the previous case, $F(x)$ belongs to the set $\bar{x}+tv_{j}$, where $t\in[t(y)-d(x,y),t(y)+d(x,y)]$. Since $d(o_{1},y)\leq d(o_{1},x)$, the point in this set furthest from $F(y)$ is $\bar{x}+(t(y)+d(x,y))v_{j}$. Denote the latter point by $\widetilde{x}$. Let $\widetilde{y}=-|\bar{y}|u_{1}+t(y)v_{1}$, that is, the point along the negative horizontal axis at the same distance from the origin as $\bar{y}$. The point $\widetilde{y}$ belongs to the line $\{\bar{y}+tv_{1}:t\in\mathbb{R}\}$ but is further away from $\widetilde{x}$ than $F(y)$. Thus $\|F(x)-F(y)\|\leq\|\widetilde{x}-\widetilde{y}\|$. But $\|\widetilde{x}-\widetilde{y}\|\leq\sqrt{3}d(x,y)$ by the same analysis as in 3.1. This establishes the upper bound. ∎
$F(y)$$F(x)$$o$

### v2

@@ Abstract.
footnotetext: 2020 Mathematics Subject Classification. 51F30, 30L05.
1. Introduction
The bi-Lipschitz embedding problem asks one to give an intrinsic characterization of those metric spaces that can be embedded in some Euclidean space under a bi-Lipschitz map. Moreover, if a metric space does embed in some Euclidean space, one may then attempt to find or estimate the optimal target dimension and bi-Lipschitz distortion. The bi-Lipschitz embedding problem is wide open in full generality and has produced a rich body of research from a variety of mathematical points of view. We refer the reader to [6], [7], [10, Chapter 15], [11], [13] and [14] for general background.
In this article, we study a specific class of metric spaces, namely, metric polygons, and their embeddings in Euclidean spaces, with a focus on metric triangles. We say that a metric polygon is a metric space comprised of finitely many closed intervals (i.e., metric spaces each isometric to $[0,a]$ for some $a>0$), called edges, joined in cyclic fashion. For two points lying on different edges, there is no constraint on the distance between them beyond what follows from the triangle inequality. A metric polygon with three edges is called a metric triangle, with four edges is called a metric quadrilateral, and so forth. See ˜2.4 below for a formal definition.
To our knowledge, metric polygons first appear explicitly in the recent paper [12], in which it is shown that any metric triangle embeds in $\mathbb{R}^{2}$ (equipped with the Euclidean metric) under a bi-Lipschitz map with distortion at most $12$. The fact is then applied to the problem of finding good polyhedral approximations of metric surfaces under minimal geometric assumptions. We refer to the embedding found in [12] as the tripodal embedding; it will be reviewed in Section˜3 below. Our goal is to explore the topic of embedding metric polygons more systematically and extend the embedding result from [12] in several directions.
@@ 1. Introduction
In Section˜4.4, we use an example that we call the twisted heart to show that the constant $4\sqrt{7/3}$ cannot be improved. To our knowledge, this example has not been considered in the literature previously. This leaves open the question of the best possible value for the distortion $D$ with an arbitrary embedding. Let $\operatorname{lip}(\triangle)$ denote the minimum value for which any metric triangle embeds in $\mathbb{R}^{2}$ with distortion at most $\operatorname{lip}(\triangle)$. A simple example, called the three-petal rose, shows that $\operatorname{lip}(\triangle)\geq 2$; see Section˜4.2. The example of the twisted heart suggests an even larger value for $\operatorname{lip}(\triangle)$, perhaps roughly the value $4$. However, we have not been able to find any general embedding method that improves upon the tripodal embedding.
@@
In Section˜4, we study the previously mentioned examples, along with two other representative examples: tripods and the intrinsic circle. Specifically, we consider the problem of finding distortion-minimizing bi-Lipschitz embeddings for these triangles. We can think of such a distortion-minimizing embedding as an optimal “Euclidean realization” of the original triangle. Note, however, that such an embedding is typically not unique; since the distortion of an embedding depends only on the worst pair of points in the space, one can typically perturb a given embedding elsewhere without affecting the distortion.
In the final main section of this paper, Section˜5, we consider the situation for arbitrary metric polygons. Note that it follows immediately from general properties of bi-Lipschitz embeddings (see Theorem 3.2 in [7]) that every metric $n$-gon bi-Lipschitz embeds in $\mathbb{R}^{2n+1}$ with uniformly bounded distortion depending only on $n$. Thus our main question is, for a given $n$, to determine the smallest dimension for which one obtains an embedding of every $n$-gon with uniformly bounded distortion. We show that the analogue of ˜1.1 is false for metric quadrilaterals, even with simple topology.
@@
There is a sequence of metric quadrilaterals $X_{n}$, each homeomorphic to the circle, such that $\operatorname{lip}(X_{n},\mathbb{R}^{2})$ is unbounded as $n\to\infty$.
@@
We also ask whether every metric quadrilateral is individually bi-Lipschitz embeddable in $\mathbb{R}^{2}$, i.e., without a uniform bound on the bi-Lipschitz constant.
@@
On the other hand, $K_{5}$, the complete graph on five vertices, can be realized as a metric pentagon. That is, a metric pentagon exists that is homeomorphic to $K_{5}$. See ˜5.3 below. We recall Kuratowsk’s Theorem in graph theory, which states that a graph can be topologically embedded in the plane if and only if it does not contain a subgraph that is a subdivision of $K_{5}$ or $K_{3,3}$ (the complete bipartite graph on two sets of three vertices). Thus a metric pentagon need not be even topologically embeddable in the plane.
In addition to our work, there have been other investigations into the metric geometry of classes of one-dimensional metric spaces. For example, low-distortion embeddings of metric graphs in Euclidean space were studied by Linial–London–Rabinovich in [8]. More recently, bi-Lipschitz embeddings of so-called quasiconformal trees have also been studied by David–Eriksson-Bique–Vellis in [5] and [4].
2. Background
2.1. Metric spaces and bi-Lipschitz maps
@@
Finally, the concatenation of two curves $\Gamma_{1}\colon[a,b]\to X$, $\Gamma_{2}\colon[c,d]\to X$ satisfying $\Gamma_{1}(b)=\Gamma_{2}(c)$ is denoted by $\Gamma_{1}*\Gamma_{2}$. This is the curve from [a,b+dc][a,b+d-c] to $X$ defined by the formula $\Gamma_{1}(t)$ if $t\leq b$ and $\Gamma_{2}(t-b+c)$ if $t\geq b$.
2.2. Metric polygons
@@
Let $n\geq 2$. A metric $n$-gon is a metric space that is the union of $n$ subspaces $I_{1},\ldots,I_{n}$ that are each isometric to a closed interval $0,\ell_{n}]$ and connected at the endpoints cyclically. More precisely, if we let $\varphi_{j}\colon[0,\ell_{j}]\to I_{j}$ denote the required isometry onto the interval $I_{j}$, then it holds that $\varphi_{j}(\ell_{j})=\varphi_{j+1}(0)$ for all $j\in\{1,\ldots,n\}$, with the identification $n+1=1$. A metric $3$-gon is called a metric triangle, and a metric $4$-gon is called a metric quadrilateral, and so forth. More generally, a metric polygon is a metric $n$-gon for some $n$.
@@
Note that this definition allows the sets $I_{j}$ to intersect at points other than the endpoints, although the nature of such additional intersection points is constrained by the triangle inequality. Also, note that the terms “edge” and “vertex” here are distinct from the topological meaning of these terms for a graph. For example, graphs (such as $K_{5}$ or $K_{3,3}$, as discussed in ˜5.3) can be represented as metric polygons. In these cases, the topological edges and vertices of the graph need not coincide with its edges and vertices as a metric polygon.
Next, we introduce some main examples of metric triangles that will be studied further in Section˜4. These are illustrated in Figure˜1. In each of these examples, the space is a length space or subset of a length space.
@@
    To view this space as a metric triangle, we pick any three points that do not all belong to the interior of the same semicircle to be vertices. For example, we can take $\theta_{1}=0$, $\theta_{2}=2\pi/3$ and $\theta_{3}=4\pi/3$.
@@
    The three-petal rose. Let $S_{1},S_{2},S_{3}$ be three copies of the intrinsic circle as just defined. Let $R$ be the wedge sum of $S_{1},S_{2},S_{3}$ with basepoint 0 in each $S_{i}$ (that is, the quotient space of the disjoint union of $S_{1},S_{2},S_{3}$ given by identifying the point 0 in each $S_{i}$). We define $d$ to be the length metric on $R$. More precisely, $d$ is defined as follows. If $\theta,\tau\in S_{i}$ for the same $i$, then $d(\theta,\tau)$ agrees with the metric in (a). If $\theta\in S_{i}$ and $\tau\in S_{j}$ for $i\neq j$, then $d(\theta,\tau)=d(\theta,0)+d(0,\tau)$.
    To view this space as a metric triangle, we take the antipodal point on each circle as vertices. That is, let $p=\pi\in S_{1}$, $q=\pi\in S_{2}$ and $r=\pi\in S_{3}$. The points 0 and $p$ divide $S_{1}$ into two semicircles $S_{1}^{1}$, $S_{1}^{2}$. Likewise, $S_{2}$ and $S_{3}$ are divided into semicircles $S_{2}^{1},S_{2}^{2}$ and $S_{3}^{1},S_{3}^{2}$, respectively. Then the edges of $R$ are $pq]=S_{1}^{2}\cup S_{2}^{1}$, $qr]=S_{2}^{2}\cup S_{3}^{1}$ and $rp]=S_{3}^{2}\cup S_{1}^{1}$.
@@
    Tripods. Choose values $l_{i}\geq 0$, $i=1,2,3$, and consider the three intervals $T_{i}=[0,l_{i}]$, $i=1,2,3$. We define $T$ to be the wedge sum of $T_{1},T_{2},T_{3}$ with basepoint 0 in each $T_{i}$. We define $d$ to be the length metric on $T$. That is, if $x,y\in T_{i}$ for the same $i$, then d(x,y)=|xy|d(x,y)=|x-y|. If $x\in T_{i}$ and $y\in T_{j}$ for $i\neq j$, then $d(x,y)=d(x,0)+d(0,y)=x+y$. Such a metric space is called a tripod, or a metric tripod for emphasis. We call the point 0 the center of the tripod.
    Let $p=l_{1}\in T_{1}$, $q=l_{2}\in T_{2}$, $r=l_{3}\in T_{3}$. We can view $T$ as a metric triangle by taking the points $p,q,r$ as vertices and $pq]=T_{1}\cup T_{2}$, $qr]=T_{2}\cup T_{3}$, $rp]=T_{3}\cup T_{1}$ as edges.
@@
    The twisted heart. The example is less obvious but turns out to be the extremal example for the tripodal embedding. We start with the graph consisting of fourteen edges of unit length as shown in Figure˜1(d). The distance between two points is the length of the shortest curve between them. We obtain a triangle by taking $p,q,r$ as vertices and removing the edges from $a$ to $c$ and from $b$ to $e$ (the red edges in Figure˜1(d)) without changing the metric on the remaining points. Observe that each pair of vertices is at distance $4$ apart. Moreover, $d(a,b)=4$ while $d(a,c)=d(b,e)=1$. We denote this example by $H$.
ppqqrr
ppqqrr
ppqqrr
ppqqrraabbeecc
Figure 1. Examples of metric polygons.
Our final definition, used in the proof of ˜1.1, is the Gromov product, which measures the deficit in the triangle inequality for a triple of points.
@@
Let $\triangle$ be a metric triangle with vertices $p,q,r$. Let $T$ be the tripod formed as in ˜2.5 by taking $l_{1}=(q\cdot r)_{p}$, $l_{2}=(p\cdot r)_{q}$ and $l_{3}=(p\cdot q)_{r}$. Define a map $P\colon\triangle\to T$ by setting $P(p)=l_{1}\in T_{1}$, $P(q)=l_{2}\in T_{2}$, $P(r)=l_{3}\in T_{3}$, and mapping each edge of $\triangle$ by arc length onto the corresponding edge of $T$. Then $P$ is $1$-Lipschitz.
@@
3. Proof of optimal distortion bound
In this section, we give the proof of our main result, ˜1.1, giving the optimal distortion bound for the tripodal embedding of an arbitrary metric triangle.
We start by reviewing the definition of the tripodal embedding as given in [12]. Let $\triangle$ be an arbitrary metric triangle with edges $I_{1},I_{2},I_{3}$. Denote the metric on $\triangle$ by $d$. Let $p$ be the common vertex of $I_{1}$ and $I_{3}$, $q$ be the common vertex of $I_{1}$ and $I_{2}$, and $r$ be the common vertex of $I_{2}$ and $I_{3}$. For each edge $I_{j}$, let $\widehat{I}_{j}$ denote the union of the other two edges of $\triangle$. We can associate to the triangle $\triangle$ a metric tripod $T$ with matching edge lengths and corresponding map $P\colon\triangle\to T$ as in ˜2.7. To this $T$ we associate an embedded tripod $Y$ and map $\varphi\colon T\to Y$ following the notation of the previous section. For conciseness, we write $\bar{x}$ in place of $(\varphi\circ P)(x)$. Let $v_{1}=(\frac{1}{2},\frac{\sqrt{3}}{2})$, $v_{2}=(-1,0)$ and $v_{3}=(\frac{1}{2},-\frac{\sqrt{3}}{2})$. Then $F\colon\triangle\to\mathbb{R}^{2}$ is defined by the formula
@@ 3. Proof of optimal distortion bound
[…]
Figure 2. The five cases in the proof of ˜1.1.
@@
We fix some additional notation. Let $I_{1}^{p}$ denote the subarc of $I_{1}$ consisting of those points $x\in I_{1}$ for which $d(x,p)\leq(q\cdot r)_{p}$. Define $I_{1}^{q}$, $I_{2}^{q}$, $I_{2}^{r}$, $I_{3}^{r}$ and $I_{3}^{p}$ similarly. For each edge $I_{j}$, let $o_{j}$ denote the unique point such that $\bar{o}_{j}$ is the origin. Observe that $I_{1}=I_{1}^{p}\cup I_{1}^{q}$, with $I_{1}^{p}$ and $I_{1}^{q}$ intersecting at the point $o_{1}$, and likewise for $I_{2}$ and $I_{3}$.
In this proof, we consider two arbitrary distinct points $x,y\in\triangle$ and show that $\|F(x)-F(y)\|$ satisfies appropriate bounds. By symmetry, we may consider five individual cases in our proof of ˜1.1: (1) $x,y\in I_{1}^{p}$; (2) $x\in I_{1}^{p},y\in I_{1}^{q}$; (3) $x\in I_{1}^{p},y\in I_{3}^{p}$; (4) $x\in I_{1}^{p},y\in I_{3}^{r}$; (5) $x\in I_{1}^{q},y\in I_{3}^{r}$. See Figure˜2. Each of these cases is represented by one of the lemmas in this section. Next, given this choice of $x$ and $y$, we let $z\in\triangle$ denote a point satisfying $d(x,z)=\operatorname{dist}(x,\widehat{I}_{1})$. This point may fail to be uniquely determined, so we fix a choice of $z$. Similarly, we let $w\in\triangle$ denote a point satisfying $d(y,w)=\operatorname{dist}(y,\widehat{I}_{j})$, where $j$ is such that $y\in I_{j}$.
@@
Assume that $d(x,p)<d(y,p)$. We first verify the lower bound. Note that both $\bar{x}$ and $\bar{y}$ lie on the positive real axis, with $\|\bar{x}-\bar{y}\|=d(x,y)$. Moreover, $F(x)$ and $F(y)$ lie on the lines $\{\bar{x}+tv_{1}:t\in\mathbb{R}\}$ and $\{\bar{y}+tv_{1}:t\in\mathbb{R}\}$, respectively. The distance between these lines is $\frac{\sqrt{3}}{2}d(x,y)$. See Figure˜3(a). This verifies the lower bound.
Next, we show the upper bound. Observe that $|d(x,\widehat{I}_{1})-d(y,\widehat{I}_{2})|\leq d(x,y)$. Write $F(y)$ as $F(y)=\bar{y}+t(y)v_{1}$ for some $t(y)\geq 0$. Then $F(x)$ belongs to the set $\bar{x}+tv_{j}$, where $t\in[t(y)-d(x,y),t(y)+d(x,y)]$. The point in this set furthest from $F(y)$ is $\bar{x}+(t(y)+d(x,y))v_{j}$, which is at distance $\sqrt{3}d(x,y)$ from $F(y)$. See Figure˜3(b). This verifies the upper bound. ∎
@@
Assume without loss of generality that $d(o_{1},y)\leq d(o_{1},x)$. We first verify the lower bound. As in the previous case, observe that $F(x)$ and $F(y)$ lie on the lines $\{\bar{x}+tv_{1}:t\in\mathbb{R}\}$ and $\{\bar{y}+tv_{1}:t\in\mathbb{R}\}$, respectively. The distance between these lines is $\frac{\sqrt{3}}{2}d(o_{1},x)+\frac{\sqrt{3}}{2}d(o_{1},y)$. It follows that
$\frac{\sqrt{3}}{2}(d(o_{1},x)+d(o_{1},y))\leq\|F(x)-F(y)\|.$
See Figure˜4(a). Since $d(x,y)=d(x,o_{1})+d(o_{1},y)$, we obtain the lower bound.
We next verify the upper bound. Write $F(y)$ as $F(y)=\bar{y}+t(y)v_{1}$ for some $t(y)\geq 0$. As in the previous case, $F(x)$ belongs to the set $\bar{x}+tv_{1}$, where $t\in[t(y)-d(x,y),t(y)+d(x,y)]$. Since $d(o_{1},y)\leq d(o_{1},x)$, the point $\bar{x}+(t(y)+d(x,y))v_{1}$ has largest possible distance from $F(y)$ in this set. Denote the latter point by $\widetilde{x}$. Let $\widetilde{y}=(-d(o_{1},y),0)+t(y)v_{1}$; that is, in the definition of $F(y)$, we replace $\bar{y}$ with the point along the negative horizontal axis at the same distance from the origin. The point $\widetilde{y}$ belongs to the same line $\{\bar{y}+tv_{1}:t\in\mathbb{R}\}$ but is further away from $\widetilde{x}$ than $F(y)$. Thus $\|F(x)-F(y)\|\leq\|\widetilde{x}-\widetilde{y}\|$. But $\|\widetilde{x}-\widetilde{y}\|\leq\sqrt{3}d(x,y)$ by the same analysis as in ˜3.1. This establishes the upper bound. ∎
F(y)F(y)F(x)F(x)$o$